If it takes 7 cups of flour to make 4 loaves of bread, how many loaves of bread can you make with 35 cups of flour?
Answer by checkley71(8403)

7/4=35/X CROSS MULTIPLY
7X=4*35
7X=140
X=140/7
X=20 LOAVES CAN BE MADE WITH 35 CUPS OF FLOUR.